OpenStudy (simplymarie_x):

What is the mass, in grams, of a sample of 6.98 × 1024 atoms of magnesium (Mg)? Show your work or explain the steps that you used to determine your answer.

OpenStudy (anonymous):

6.98x10^24 atoms = n mols = n*6.022x10^23 n= (69.8x10)/(6.022x10)= (698)/(60.22) = 11.6 mol mg 24.305(molar mass of Mg)g/mol = Xg/11.6 mol mg (24.305 g/mol)/11.6 mol mg = X (grams) 2.09 g magnesium